OpenStudy (anonymous):

Evaluate log base8 12.

OpenStudy (saifoo.khan):

Try this one!

OpenStudy (turingtest):

yes, do please try... we will guide you from there

OpenStudy (anonymous):

ok what do i do first

OpenStudy (turingtest):

change of base as was mentioned by saifoo and others in your last post

OpenStudy (anonymous):

the base of what?

OpenStudy (turingtest):

looks like you want to simplify though??? what is the question here?

OpenStudy (anonymous):

im so confused

OpenStudy (anonymous):

lemme give u a simple method log base 2 (16) = x can be written as 2^x =16 try to figure out the pattern here from previous problems

OpenStudy (turingtest):

if you want to use a calculator you probably need to change the base of the logarithm if you are trying to simplify you can apply the basic log property\[\log(ab)=\log a+\log b\] @niravshah08 has shown a good way to look at the problem

OpenStudy (anonymous):

...lost, just.... lost

OpenStudy (saifoo.khan):

Change of base is something as \[\Large \log_a b \to \frac{ \log b}{\log a}\]

OpenStudy (anonymous):

mabey this would be a good time to mention that i dont know how to use my calculator

OpenStudy (anonymous):

u dont need calculator for these problems

OpenStudy (saifoo.khan):

which calc do u have? @nyan21

OpenStudy (lgbasallote):

@saifoo.khan just gave you a hint on how to use a calculator on these kinds of problems

OpenStudy (turingtest):

you need to understand what a logarithm is... but you just told me two posts ago that you are dyslexic and "can't do math" so I recommend you go to your doctor and get a note excusing you from math class I've seen it done before, it's possible if you think you can learn it though I'll be happy to teach you the idea of a logarithm

OpenStudy (anonymous):

log base8 12 can be evaluated as 8^x = 12 do u see any connection ?

OpenStudy (lgbasallote):

\(\log_8 12\) is like asking..what number will you raise 8 to to get 12

OpenStudy (turingtest):

^exactly the explanation I use to describe what a logarithm means focus on what @lgbasallote has said above

OpenStudy (anonymous):

i give up

OpenStudy (turingtest):

\[\log_24\]is like asking "what power do you raise 2 to in order to get 4?" what is the answer to that question?

OpenStudy (anonymous):

i think my brain just exploded

OpenStudy (anonymous):

im trying to understand this but its like your speaking japanese

OpenStudy (turingtest):

well when your brain reconstructs itself let us know, and we will help you understand the idea of the logarithm \[\log_24\]is asking "2 raised to what power equals 4?" please try to meditate on this question; that will help you understand the concept of the logarithm.

OpenStudy (anonymous):

2?

OpenStudy (turingtest):

yes :) now what is\[\log_2(8)\]?

OpenStudy (anonymous):

3?

OpenStudy (anonymous):

u gpt it :)

OpenStudy (turingtest):

very good! so do you think you are beginning to understand the idea of the logarithm?

OpenStudy (anonymous):

no still dont get it but 3 isnt in my mutiple coice answers i have 0.84 0.67 1.50 1.19

OpenStudy (turingtest):

then you are supposed to use a calculator... what kind do you have

OpenStudy (anonymous):

literally it just died and idk it doesnt say

OpenStudy (turingtest):

well then use wolfram alpha for this kind of thing http://www.wolframalpha.com/ post your question as log_8(12) or something and it will give you the answer

OpenStudy (turingtest):

of course you would be much better off using the change of base formula and the calculator in your computer

OpenStudy (anonymous):

thankx

OpenStudy (turingtest):

welcome if you want to use that site as a crutch you \(will\) fail your exams though. just to let you know...

OpenStudy (anonymous):

well im doing this onlin eclass so i dont really need to know this stuff i just use the internet for answers